If I=V/R and V=250 volts and R=50ohms,find the change in I resulting from an increase of 2 volts in v and an increase of 0.5ohm in R. Using total differential method.

1
Expert's answer
2022-05-03T16:27:59-0400

The total differential for I is

"dI=\frac{1}{R}dV-\frac{V}{R^2}dR"

Approximations: given small values for "\Delta V" and "\Delta R" :

"\Delta I=\frac{1}{R}\Delta V-\frac{V}{R^2}\Delta R=\frac{2}{50}-\frac{250\cdot 0.5}{50^2}=-0.01"

Answer: "\Delta I=-0.01" A.